i was studying for the ACT compass test and i am having problems with functions. the question was: if f(4)=0 and f(6)=6 then which of the following could represent f(x)? The answer is 3x-12, but i do not understand how to get to that answer, what does the 4 and 6 have to do with it? thank you ahead of time
Found 2 solutions by jim_thompson5910, Edwin McCravy:
Answer by jim_thompson5910(35256)

Hint: If f(4)=0, then we have the point (4,0). Recall that f(x) is really 'y'. So if f(4)=0, then x=4 and y=0. Similarly, if f(6)=6, then we have the second point (6,6)

Answer by Edwin McCravy(20059)


Do you know what f(4) and f(6) means?  Apparently not.

f(4) means what you get when you substitute (4) for x in the right side of 

f(x) = 3x-12

If you put a (4) wherever the x is you get

f(4) = 3(4)-12 and if you simplify the right side you get

f(4) = 12-12

or

f(4) = 0

---

Similarly, f(6) means what you get when you plug 6 in for x in
the right side of 

f(x) = 3x-12

If you put a (6) wherever the x is you get

f(6) = 3(6)-12 and if you simplify the right side you get

f(6) = 18-12

or

f(6) = 6

--------------------

Now if you tried doing those two substitutions in any of the other

choices you had on your test you would not get 0 for f(4) and also 

6 for f(6).

Learn that "f( )" with a number in the parentheses means 

to substitute whatever number is in the parentheses for x into

the right side of the equation which has "f(x) = ____" where the 

right side is some algebraic expression that contains x, such as

3x-12 or 6x-3 or 7x+3 or 8x²+3x-9 or etc. etc. etc.  Then after

substituting, then  simplify, and the number you get is equal 

to f(that number)

Edwin